If I had an old TV that is 15" by 20" and bought a new TV that measures 18" by 32", how do I make a formula to find out the percent of increase in area of the new TV Screen? I think I can find area of each, but am lost when it comes to finding the percent of increase. Can you help please
Answer by JimboP1977(311)

Original area = 15 * 20 = 300 inches^2
New area = 18 * 32 = 576 inches^2
To find the percentage increase we take the difference of the two which is 276.
Then we take the difference and divide by the original area:
276/300 = 0.92
Multiply by 100 to give a percentage = 92% increase.
To think about it another way we are saying that the area has increased by 92% of its original size.
So 92% of 300 is 276 which is the amount of the increase.
Does that make any sense at all?
Do you need further clarification?